OpenStudy (anonymous):

square root 16x + square root x^5

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

\[\sqrt{16 x}+\sqrt{x^5}=4\sqrt{x}+\sqrt{x^4*x}=4\sqrt{x}+x^2\sqrt{x}=\sqrt{x}\left( 4+x^2 \right)\]